2. A telecommunication company is investing $65,000 in a signal enhancing system. The system salvage value is 0 and MARR is ePlease do not use Excel
Answer #1

Initial Cost = 65,000

The savings are

Year 1 = 25,000

Year 2 = 30,000

Year 3 = 30,000

Year 4 = 40,000

Year 5 = 46,000

MARR = 18%

1. Future Worth of the Project

FW = -65,000 (F/P, 18%, 5) + 25,000 (F/P, 18%, 4) + 30,000 (F/P, 18%, 3) + 30,000 (F/P, 18%, 2) + + 40,000 (F/P, 18%, 1) + 46,000

FW = -65,000 (2.2878) + 25,000 (1.9388) + 30,000 (1.6430) + 30,000 (1.3924) + 40,000 (1.1800) + 46,000

FW = 84,025

The FW is greater than zero, the project is a good project.

2. Calculating the IRR using the trial and error method.

Let the rate of interest is 38%. Calculate the PW of the cash flows at 38%.

EOY

Cash Flows

Discounting Factor at 38%

Present Value

0

-65000

1.00000

-65000

1

25000

0.72464

18115.94

2

30000

0.52510

15752.99

3

30000

0.38051

11415.21

4

40000

0.27573

11029.19

5

46000

0.19980

9190.992

504.3304

The PW is 504.33 and it’s a positive PW. So, increase the rate of interest to get a negative PW. Increase the rate of interest to 39% and calculate PW.

EOY

Cash Flows

Discounting Factor at 39%

Present Value

0

-65000

1.00000

-65000

1

25000

0.71942

17985.61

2

30000

0.51757

15527.15

3

30000

0.37235

11170.61

4

40000

0.26788

10715.21

5

46000

0.19272

8865.104

-736.316

The PW is -736.32

Using interpolation

IRR = 38% + [504.33 – 0 ÷ 504.33 – (-736.32)] * 1%

IRR = 38.40%

The IRR is 38.40%

3. Discounted Payback Period

Year

CF

PV Factor

DCF

CCF

0

$-65,000

1

$-65,000

$-65,000

1

$25,000

0.85

$21,186.44

$-43,813.56

2

$30,000

0.72

$21,545.53

$-22,268.03

3

$30,000

0.61

$18,258.93

$-4,009.1

4

$40,000

0.52

$20,631.56

$16,622.45

5

$46,000

0.44

$20,107.02

$36,729.48

Payback Period = 3 + [-4009.1 – 0 ÷ 4009.1 – (16,622.45)] * 1 = 3.16 years

4. Simple Payback Period

Year

CF

NCF

0

$-65,000

$-65,000

1

$25,000

$-40,000

2

$30,000

$-10,000

3

$30,000

$20,000

4

$40,000

$60,000

5

$46,000

$106,000

Payback Period = 2 + (10,000 ÷ 30,000) = 2.33 years